A spherical ball of surface area 20 cm2 absorbs any radiation that falls on it. It is suspended in a closed box maintained at 57°C. (a) Find the amount of radiation falling on the ball per second. Sol. (a) A = 20 cm^2 = 20× 10^–4 m^2, T = 57°C = 330 K E = A σT^4 = 20 × 10^–4 × 6 × 10^–8 × (330)^4 × 10^4 = 1.42 J COULD U HELP ME FIGURE OUT FROM WHERE THE LAST 10^4 CAME FROM?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
Sol. (a) A = 20 cm^2 = 20× 10^–4 m^2, T = 57°C = 330 K E = A σT^4 = 20 × 10^–4 × 6 × 10^–8 × (330)^4 × 10^4 = 1.42 J 
Hope it helps
Regards
Arun (askIITians forum expert)